OpenStudy (anonymous):

3sinx = 2cos^2x solve.

OpenStudy (mertsj):

\[3\sin x=2(1-\sin ^2x)\] \[2\sin ^2x+3\sin x-2=0\] \[(2\sin x-1)(\sin x+2)=0\] You can finish it now.

OpenStudy (anonymous):

Thank you, i got up to the step before the answer you gave me...

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

So what did you get for x?

OpenStudy (anonymous):

are there 2 answers for x?

OpenStudy (anonymous):

Sin x = 1/2

OpenStudy (anonymous):

cause it is also [0,2pi)

OpenStudy (mertsj):

If sinx = 1/2, then x = pi/6 and 5pi/6

OpenStudy (anonymous):

ok, thanks

OpenStudy (mertsj):

yw